what is the length of the rectangular table if the the length is 18 inches longer than twice the width if the perimeter is 252
Answer by Cromlix(4381)

Width = x
Length = 18 + 2x
Perimeter = 2* width + 2 * length
252 = 2(18 + 2x) + 2(x)
252 = 36 + 4x + 2x
252 = 36 + 6x
252 - 36 = 6x
6x = 216
x = 36
Width = 36 inches
Length = 90 inches.
Hope this helps.
:-)